# Latch, Spring and Release: The Efficiency of Power-Amplified Jumping

### v1

@@ Abstract
Figure 4: Energy efficiency of jumping of a curved shell given by the ratio of translational kinetic energy at $T=T_{\text{off}}$ to stored energy at $T=0$ as a function of the latch lifting time scale. Theoretical predictions from linearized theory are depicted in solid lines. Scaled values of the efficiency for L. biguttatus [7], jumping hoops [5], and our experiments of a curved strip are also included as symbols. Solid lines for $\beta=1$ and dashed lines for $\beta=10$. Colors range from dark brown (small $H$) to yellow (large $H$) for $H=0.1,1,2,10$,
@@
We have developed a dynamical theory that describes the observed efficiency of the jumps of insect larvae, as well as providing a mechanism for the wide distribution of observed jump efficiencies: an exponential decay means that small changes in $\tau$ lead to significantly less efficient jumps. While this may pose a substantial challenge to insect larvae, it also focuses the design criteria for controlling jumps in robots [6], for example. To make this a reality, however, may require generalizing this work to scenarios where the contact set is more complicated than that considered here — for example, in insect-inspired jumping robots, the contact region grows during the early stages of jumping, before ultimately decreasing [6].
Acknowledgements.
@@
This work was partially supported by the UK Engineering and Physical Sciences Research Council via Grant No. EP/W016249/1 and Grant No. EP/Y027949/1 (M.S. and D.V.), and by the Swedish Research Council under Grant No. 638-2013-9243 (M.S. and J.W.). For the purpose of Open Access, the authors will apply a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.

### v2

@@ Abstract
Figure 4: Energy efficiency of jumping of a curved shell given by the ratio of translational kinetic energy at $T=T_{\text{off}}$ to stored energy at $T=0$ as a function of the latch lifting time scale. Theoretical predictions from linearized theory are depicted in solid lines. Scaled values of the efficiency for L. biguttatus [12], jumping hoops [5], and our experiments of a curved strip are also included as symbols. Solid lines for $\beta=1$ and dashed lines for $\beta=10$. Colors range from dark brown (small $H$) to yellow (large $H$) for $H=0.1,1,2,10$.
Our results for $\mathcal{E}(\tau)$ are shown for a variety of values of $\beta$ and $H$ in Fig. 4. As might have been expected, these show that the efficiency of jumping decreases with $\tau$, though it is somewhat surprising that our results suggest that this decay is exponential. Moreover, there appear to be certain pessimal values of $\tau$ where the efficiency drops even more dramatically. We attribute this to a destructive interference between the imposed decay of $F(t)$ and the natural relaxation time of the shell, which is consistent with the re-entrant efficiency observed for lighter, stiffer shells. (Using this idea, we estimate these pessimal values in the SI using asymptotic methods.) Importantly, these pessimal efficiencies arise from the interplay between the latch-release dynamics and the elastic deformation of the jumping shell.
This exponential decay of the efficiency is not a mere consequence of the exponential force release (3). Indeed, we show (see SI Appendix, Fig. S11B) that different force protocols in a simplified 1D mass-spring model (following Ref. [27]) give efficiency of jumping decays markedly slower than exponential. This result underscores the importance of integrating force control of latch release with a geometric model of the jumper’s elastodynamic behavior. The sensitive dependence of the efficiency of jumping on the release time scale $\tau$ that we observe suggests that the dynamic deformation of the strip itself accentuates this – a feature that is not captured in the LaMSA formalism [9, 10]. Moreover, this sensitivity suggests a rationalization for the wide variability in the efficiency of jumps that were reported by [12]: even an apparently small change in the latch release time $\tau$ can lead to a large decrease in the efficiency of jumping.
@@
We have developed a dynamical theory that describes the observed efficiency of the jumps of insect larvae, as well as providing a mechanism for the wide distribution of observed jump efficiencies: an exponential decay means that small changes in the time scale of latch release, $\tau$, can lead to significantly less efficient jumps. This may pose both a challenge and an opportunity to insect larvae: while the timing of release is important, it also gives a new means of modifying the jump after the energy has been stored and the spring ‘primed’. This may also help to focus the design criteria for controlling jumps in robots [6], for example, providing another setting of insect-inspired robotics [28].
Our geometrical theory integrates latch-mediated energy release and useful work generation to amplify power, thereby providing a geometrical description of jumping that unifies the most relevant aspects of LaMSA systems [9, 10, 11] and jumping robots [5, 18, 6, 19, 20].
@@
This work was partially supported by the UK Engineering and Physical Sciences Research Council via Grant No. EP/W016249/1 and Grant No. EP/Y027949/1 (M.S. and D.V.), and by the Swedish Research Council under Grant No. 638-2013-9243 (M.S. and J.S.W.). For the purpose of Open Access, the authors will apply a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission. M.S. thanks Maria Bruna for helpful comments.